OpenStudy (anonymous):

Can someone please help me find the approximation of cubed root of 1002

OpenStudy (goformit100):

USE DIFFRENTIAL CALCULUS

OpenStudy (anonymous):

Yeah I know what I have to use but I still don't get it.

OpenStudy (goformit100):

WHAT PROBLEM ARE YOU FACING ?

OpenStudy (anonymous):

I have fx=x^1/3 and f'x= 1/3 x^-2/3 and thats all I know. And that cubed root of 1000 is 10 and my h is 2 i dont know how to use that info

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):